OpenStudy (anonymous):

simplify: (2x - 7)^2

OpenStudy (anonymous):

4x^2-28x+49

OpenStudy (anonymous):

\[4x^2-28x+49\]

OpenStudy (diyadiya):

\[(a-b)^2=a^2-2ab+b^2\]\[(2x-7)^2=(2x)^2-2*2x*7+7^2\]\[=4x^2-28x_49\]

OpenStudy (diyadiya):

\[4x^2-28x + 49\]